graph an equation of the form Ax + By = C

3x - 2y = 8

Answer by shantysaggu007(3)

graph an equation of the form Ax + By = C
3x - 2y = 8
Compare Ax=3
bX=-2
c=8
L.HS=3X-2Y…………………..EQ (1)
PUT VALUE X=1,Y=1 IN EQ ….(1)
3(1)-2(1)=8
3-2=8
1=8
L.H.S#R.H.S (1,1)
PUT VALUE X=2,Y=2 IN EQ….(1)
3(2)-2(2)=8
6-4=8
2=8
L.H.S#R.H.S (2,2)
PUT VALUE X=4,Y=2 IN EQ….(1)
3(4)-2(2)=8
12-4=8
8=8
L.H.S=R.H.S (4,2)
MAKE GRAPH ACCORDING TO PUTPUT (FIRST X, SECOD Y)